OpenStudy (anonymous):

(2x^2 - 2) how would you find the zero of this??? Would you add 2 to the 0 and then square root it and then divide 2???

OpenStudy (anonymous):

2x^2-2=0?

OpenStudy (anonymous):

to find the zero of a polynomial im confused XD

OpenStudy (anonymous):

Order of operations. You'd have to divide by 2 first, then take the square root

OpenStudy (anonymous):

yes make it x^2-1 then you have your answer

OpenStudy (anonymous):

theres 2 answers

OpenStudy (anonymous):

So once I subtract 2, divide it by 2, and then square root it?

OpenStudy (anonymous):

x^2-1=0 ->x^2=1

OpenStudy (anonymous):

im so confused now

OpenStudy (anonymous):

oooo i got it now

OpenStudy (anonymous):

divide the whole equation by 2 then do the order of operations thanks guys

OpenStudy (anonymous):

yw